OpenStudy (anonymous):

acceleration due to gravity is g ms^-2. find the effective value of acceleration due to gravity at a height of 32km from sea level

OpenStudy (jamesj):

Again, use F = - GMm/r^2

OpenStudy (anonymous):

i used not getting ans that is why i m posting all ques i posted here i already tried them sir i want to see the calculus or u guide me i ll see where i m making mistake

OpenStudy (anonymous):

u think that we have to use thi formula GM/R^2=GR^2/(R+H)^2

OpenStudy (anonymous):

Mass of the Earth is about 5.97 * 10^24 kg, radius about 6378 km. \[\Large \begin{array}{l} g = G\frac{M}{{{r^2}}}\\ = G\frac{{5.97 \times {{10}^{24}}}}{{{{((6378 + 32) \times {{10}^3})}^2}}} \end{array}\]

OpenStudy (anonymous):